I don't understand why x=0.0033. When I put 0.0033 to the 2400=(x)(x) / (6.8*10^-3 -2x), it doesn't work. I get 0.33 for x. What did I do wrong?

5. The reaction 2 NO (g) 2 (g) O2 (g) has a value  
Answer #1

Dear yes you are wrong,

Actual answer is X = 0.00339 as per below

2400=(x)(x) / {(6.8*10^-3 )-2x }2

2400 = X2 / ( { (4.6 x 10-5 ) - 0.0272X + 4X2 }

0.1104 - 65.28X + 9600X2 = X2

9599X2 -65.28X + 0.1104 = 0

X2 - 0.0068X + ( 1.15 x 10-5) = 0

(X - 0.00339)2 = 0

(X - 0.00339) = 0

X = 0.00339